# Two-dimensional Brownian motion with dependent components: turning angle analysis

### v2

@@ Abstract
Brownian motion in one or more dimensions is extensively used as a stochastic process to model natural and engineering signals, as well as financial data. Most works dealing with multidimensional Brownian motion assume that each component is independent. In this article, we investigate a model of correlated Brownian motion in $\mathbb{R}^{2}$, where the individual components are not necessarily independent. We explore various statistical properties of the process under consideration, going beyond the conventional analysis of the second moment. Our particular focus lies on investigating the distribution of turning angles. This distribution reveals particularly interesting characteristics for processes with dependent components that are relevant to applications in diverse physical systems. Theoretical considerations are supported by numerical simulations and analysis of two real-world datasets: the financial data of the Dow Jones Industrial Average and the Standard and Poor’s 500, and trajectories of polystyrene beads in water. Finally, we show that the model can be readily extended to trajectories with correlations that change over time.
@@ I Introduction
Since the pioneering works of Einstein and Smoluchowski [1, 2], Brownian motion (BM) has established itself as the quintessential stochastic model for diffusive transport [3, 4, 5]. BM is now widely applied in various disciplines, including biology [6, 7], statistical physics [8], chemistry [9], quantum mechanics [10, 11], ecology [12, 13], and finance [14]. The multidimensional BM, also called the multidimensional Wiener process, extends the classical one-dimensional BM to multiple dimensions, resulting in a vector of independent or correlated components. It is a fundamental stochastic process that is widely utilized to model the dynamics of systems that involve two or more variables. Each component of a multidimensional BM follows a normal distribution with stationary and independent increments, making it an appealing tool to capture the complexity of real-world phenomena. Multidimensional BM was applied to model mobile radio channels [15], data in the aircraft industry [16], image encryption algorithms [17, 18], credit risk [19, 20], diffusion of colloids [21] and protein diffusion in lipid membranes [22, 23, 24]. Typically, Brownian motion in $d$ dimensions is defined such that its projections along each dimension are statistically independent [25]. However, environmental constraints can introduce correlations between these different projections [26, 27].
@@ II Correlated BM Model
Figure 1: Representative numerical simulations of two-dimensional Brownian motion trajectories with different correlations. In all panels $D_{1}=D_{2}=0.5$, time step $1$ and trajectory length $2^{11}$. (a) Trajectories with $\rho=0$, (b) $\rho=-0.3$ and (c) $\rho=0.6$. Each trajectory is shown with a different color (yellow, blue, or green). For clarity of presentation, trajectories were shifted by 50 (yellow) and 100 units (green) to the right. The underlying randomness is the same for all panels.
@@ III Properties of the correlated BM with spatial correlations
In this section, we present the basic properties of the two-dimensional Brownian motion with spatial correlations. First, let us analyze the dependence structure of increments $\Delta^{\delta}\mathbf{R}(t)\equiv\mathbf{R}(t+\delta)-\mathbf{R}(t)$. We will also use the notation $\Delta^{1}\equiv\Delta$ for increments over a unit time. Let us first consider the dependence structure on the same coordinate. The autocovariance function of such increments, closely related to the velocity autocorrelation function (VACF)[48], is given by [49]
@@
One can easily show that such a function is symmetric in the sense $\gamma_{\Delta^{\delta}B_{1},\Delta^{\delta}B_{2}}(h)=\gamma_{\Delta^{\delta}B_{2}\Delta^{\delta}B_{1}}(h)$.Due to the construction given in eq. (2), the process is Gaussian. and the increments $\Delta^{\delta}\mathbf{R}(t)$ have a Gaussian distribution with zero mean and covariance structure given by
@@
It should be noted that, despite introducing the spatial correlation, $\rho$ does not influence MSD. The model still describes normal diffusion, just like each of its components. Figure 3 presents the MSD based on simulated 1000 trajectories of two-dimensional BM. The three considered cases ($\rho=0,\rho=-0.3$, and $\rho=0.6$ corresponding to, in order, the solid blue line, the dashed orange line, and the dashed dotted green line) exhibit linear behavior. The lines describing the dependent cases are shifted by a constant of 200 (dashed orange line) and 400 (dashed-dotted green line) for visual clarity.
@@
Figure 3: Ensemble mean square displacement of correlated BM. The solid blue line corresponds to the case with independent components ($\rho=0$), dashed orange to the negatively correlated case ($\rho=-0.3$), and the dash-dotted green line to the positively correlated case ($\rho=0.6$). The MSDs corresponding to $\rho=-0.3$ and $\rho=0.6$ are shifted by 200 and 400, respectively.
@@ III.1 Polar representation
Using the explicit form of matrix $\Sigma(t)$ given in eq. (7), we can write
@@ III.2 Turning angles
While the information given by the distribution of angles $\alpha_{t}$ is interesting, it is not necessarily useful in the case of real data, where the choice of any specific axis might be arbitrary. A more natural approach is to consider the so-called turning angles, generated by consecutive increments [52, 53, 54, 55]. In the correlated BM, the increments are still independent. This independence, as stated in the previous section, leads to the independence of the corresponding angles $\alpha_{t}^{\Delta}$. Specifically, let $\phi_{t}$ denote the angle between two consecutive increments at times $t$ and $t+1$. Figure 4 shows the angles of the increments $\alpha_{t}^{\Delta}$ and the turning angles $\phi_{t}$ in a sketched interval consisting of three consecutive observations within a trajectory. Angles $\alpha_{t}^{\Delta}$ are measured to the horizontal axis, preserving the counter-clockwise direction (cf. $\alpha_{2}^{\Delta}$ in the figure). The turning angles $\phi_{t}$ are defined by
@@
The connection to the angles $\alpha_{t}^{\Delta}$, described in the previous section, is as follows
$\phi_{t}\equiv\alpha_{t+1}^{\Delta}-\alpha_{t}^{\Delta}\ (\mathrm{mod}\ \pi).$ (14)
Since we use modulo operation, the turning angles $\phi_{t}$ are now in the interval $\left[0,\pi\right)$.
@@
For correlated BM, we can calculate the PDF of $\phi_{t}$ directly. Due to the increments of Brownian motion being independent and identically distributed (i.i.d.), we know that $\alpha_{t}^{\Delta}$’s are independent, and both $\alpha_{t}^{\Delta}$’s and $\phi_{t}$’s are identically distributed, and their distribution does not depend on time. This is not universally true and would not be the case for, for example, a two-dimensional fractional Brownian motion. Considering correlated BM, we can calculate the convolution of the PDF given in eq. (10). Let $\theta\in(-2\pi,2\pi)$. Then the density of $\alpha_{t+1}^{\Delta}-\alpha_{t}^{\Delta}$, prior to the modulo operation, is given by
$g_{\phi}(\theta)$ $=(p_{\alpha}*\overline{p}_{-\alpha})(\theta)$
$=\frac{1-\rho^{2}}{4\pi^{2}}\int_{\max\{0,\theta\}}^{\min\{2\pi,2\pi+\theta\}}\frac{1}{\cos^{2}(\theta)\sqrt{\frac{D_{2}}{D_{1}}}+\sin^{2}(\theta)\sqrt{\frac{D_{1}}{D_{2}}}-2\rho\sin(\theta)\cos(\theta)}\times$
@@
$g_{\phi}(\theta)$ $=(p_{\alpha}*\overline{p}_{-\alpha})(\theta)=\frac{1-\rho^{2}}{4\pi^{2}}\int_{\max\{0,\theta\}}^{\min\{2\pi,2\pi+\theta\}}\frac{1}{1-\rho\sin(2x)}\frac{1}{1+\rho\sin(2(\theta-x))}\mathrm{d}x,$ (16)
@@
Taking into consideration the modulo operation, we arrive at the final formula for the density of $\phi_{t}$,
$p_{\phi}(\theta)=g_{\phi}(\theta)+g_{\phi}(\theta+\pi)+g_{\phi}(\theta-\pi)+g_{\phi}(\theta-2\pi).$ (17)
Let us consider one more specific case, namely, uncorrelated coordinates with the same variance, i.e. the isotropic case. In such a case, we have
@@
$g_{\phi}(\theta)=\begin{cases}\frac{1}{(2\pi)^{2}}\theta+\frac{1}{2\pi},\quad\theta\in(-2\pi,0),\\ \frac{-1}{(2\pi)^{2}}\theta+\frac{1}{2\pi},\quad\theta\in(0,2\pi).\end{cases}$ (18)
Having the formula for $g_{\phi}$ in eq. (18), we can plug it into eq. (17) to obtain the PDF after modulo operation. Indeed, for $\theta\in(0,\pi)$ we have
$p_{\phi}(\theta)$ $=g_{\phi}(\theta)$ $+g_{\phi}(\theta+\pi)$ $+g_{\phi}(\theta-\pi)$ $+g_{\phi}(\theta-2\pi)$
@@
For comparison, we consider $D_{1}=2,D_{2}=1$ and $\rho=0.7$. The PDF of the turning angles is presented in Figure 5. We can see a clear deviation from the uniform distribution. The solid blue line corresponds to the estimated histogram of the turning angles from the simulated data (based on 1000 trajectories of length $2^{10}$ with a unit time step), while the dashed yellow line corresponds to eq. (17).
@@
Figure 5: Distribution of turning angles $\phi_{t}$ for the process with $D_{1}=2$, $D_{2}=1$, and $\rho=0.7$. The solid blue histogram corresponds to the angles calculated using eq. (13) from simulated data, and the dashed yellow line corresponds to eq. (17).
@@ IV Data
In this section, we present the results for two vastly different datasets. The first is a financial dataset that encompasses two of the main US market indices, the Dow Jones Industrial Average (DJIA) and the Standard and Poor’s 500 (S&P 500), for which we analyze their logarithmic returns. The second one relates to a soft-matter physical system, namely the motion of polystyrene beads in water. Despite obtaining universal results for any variances $D_{1},D_{2}$ related to different directions, we choose first to normalize each of the considered coordinates by their corresponding standard deviation. It is a natural procedure, especially within physical systems, where the choice of units can influence the resulting distribution of the turning angles.
@@ IV.1 Example: Financial data
In Figure 6 we present the analysis of the logarithmic returns of these two indices measured daily for a period of 2 years, from May 1, 2022, to May 1, 2024. Figure 6(a) shows the joint PDF, with the top and right histograms corresponding to DJIA and S&P500 logarithmic returns, respectively. In Figure 6(b), we present fitted Gaussian PDFs of the logarithmic returns of both indices over the selected period. The blue bars and dashed blue lines correspond to the DJIA index, while the yellow bars and dashed-dotted yellow lines correspond to S&P500. In Figure 6(c) we compare the empirical PDF of turning angles (blue bars) to the analytical density given by eq. (11).
@@ IV.2 Example: Polystyrene beads in water
The data consists of 150 two-dimensional trajectories of polystyrene beads submersed in water imaged during $40.96$ s at a frame rate of $100$ frames/s [25, 56]. The nominal bead radius is $0.6\ \mu m$. The movement of the beads can be modeled using a classical Brownian motion without correlations between the coordinates ($\rho=0$). Due to the isotropic nature of the environment, we observe that $D_{1}$ and $D_{2}$ are indeed the same. The density of increments (presented in Figure 7(a)) has the circular contour lines of the characteristic of the joint PDF of two i.i.d. random variables. The one-dimensional histograms of the data are presented in Figure 7(b) and compared with the Gaussian PDFs with matched estimated parameters ($\langle D\rangle=0.375\mu\textrm{m}^{2}$/s). In this case, given $\rho=0$, the turning angles appear to have a uniform distribution (Figure 7(c)).

### v4

@@ Abstract
Brownian motion in one or more dimensions is extensively used as a stochastic process to model natural and engineering signals, as well as financial data. Most works dealing with multidimensional Brownian motion consider the different dimensions as independent components. In this article, we investigate a model of correlated Brownian motion in $\mathbb{R}^{2}$, where the individual components are not necessarily independent. We explore various statistical properties of the process under consideration, going beyond the conventional analysis of the second moment. Our particular focus lies on investigating the distribution of turning angles. This distribution reveals particularly interesting characteristics for processes with dependent components that are relevant to applications in diverse physical systems. Theoretical considerations are supported by numerical simulations and analysis of two real-world datasets: the financial data of the Dow Jones Industrial Average and the Standard and Poor’s 500, and trajectories of polystyrene beads in water. Finally, we show that the model can be readily extended to trajectories with correlations that change over time.
@@ I Introduction
Since the pioneering works of Einstein and Smoluchowski [1, 2], Brownian motion (BM) has established itself as the quintessential stochastic model for diffusive transport [3, 4, 5]. BM is now widely applied in various disciplines, including biology [6, 7], statistical physics [8], chemistry [9], quantum mechanics [10, 11], ecology [12, 13], and finance [14]. The multidimensional BM, also called the multidimensional Wiener process, extends the classical one-dimensional BM to multiple dimensions, resulting in a vector of independent or correlated components. It is a fundamental stochastic process that is widely utilized to model the dynamics of systems that involve two or more variables. Each component of a multidimensional BM follows a normal distribution with stationary and independent increments, making it an appealing tool to capture the complexity of real-world phenomena. Multidimensional BM is ubiquitous in science and technology. To name a few examples, multidimensional BM was applied to model mobile radio channels [15], data in the aircraft industry [16], image encryption algorithms [17, 18], credit risk [19, 20], diffusion of colloids [21] and protein diffusion in lipid membranes [22, 23, 24]. Typically, Brownian motion in $d$ dimensions is defined such that its projections along each dimension are statistically independent [25]. However, environmental constraints can introduce correlations between these different projections [26, 27]. A typical example deals with financial time series, where two assets in a natural way are going in the same direction [28]. However correlated Brownian motion appears in a very wide and diverse range of systems, a non-exhaustive list includes brain networks [29], two-dimensional motion of colloidal dimers [30], optically trapped graphene flakes [31], Brownian motions on manifolds [32], and thermal motion of nanotubes [33, 34].
@@ II Correlated BM Model
Figure 1: Representative numerical simulations of two-dimensional Brownian motion trajectories with different correlations. In all panels $D_{1}=D_{2}=0.5$, time step $1$ and trajectory length $2^{11}$. (a) Trajectories with $\rho=0$, (b) $\rho=-0.3$ and (c) $\rho=0.6$. Each trajectory is shown with a different color (yellow, blue, or green). For clarity of presentation, trajectories were shifted by 50 (yellow) and 100 units (green) to the right. The underlying randomness is the same for all panels with paths with the same color corresponding to equivalent trajectories.
@@ III Properties of the correlated BM with spatial correlations
In this section, we present the basic properties of the two-dimensional Brownian motion with spatial correlations. Physical measurements are naturally sampled at discrete times, thus henceforth in the article we consider the process to be discretized at times $t=0,1,2,...$. First, let us analyze the dependence structure of increments $\Delta^{\delta}\mathbf{R}(t)\equiv\mathbf{R}(t+\delta)-\mathbf{R}(t)$. We will also use the notation $\Delta^{1}\equiv\Delta$ for increments corresponding to $\delta=1$. Let us first consider the dependence structure on the same coordinate. The autocovariance function of such increments, closely related to the velocity autocorrelation function (VACF)[55], is given by [56]
@@
Note that eqs. (4) and (6) are also valid for continuous time. One can easily show that such a function is symmetric in the sense $\gamma_{\Delta^{\delta}B_{1},\Delta^{\delta}B_{2}}(h)=\gamma_{\Delta^{\delta}B_{2}\Delta^{\delta}B_{1}}(h)$. Due to the construction given in eq. (2), the process is Gaussian and the increments $\Delta^{\delta}\mathbf{R}(t)$ have a Gaussian distribution with zero mean and covariance structure given by
@@
It should be noted that, despite introducing the spatial correlation, $\rho$ does not influence MSD. The model still describes normal diffusion, just like each of its components. Figure 3 presents the MSD based on simulated 1000 trajectories of two-dimensional BM. The three considered cases ($\rho=0,\rho=-0.3$, and $\rho=0.6$ corresponding to, in order, the solid blue line, the dashed orange line, and the dashed-dotted green line) exhibit linear behavior. The lines describing the dependent cases are shifted by a constant of 200 (dashed orange line) and 400 (dashed-dotted green line) for visual clarity.
@@
Figure 3: Ensemble mean square displacement of correlated BM. The solid blue line corresponds to the case with independent components ($\rho=0$), dashed orange to the negatively correlated case ($\rho=-0.3$), and the dash-dotted green line to the positively correlated case ($\rho=0.6$). For clarity, the MSDs corresponding to $\rho=-0.3$ and $\rho=0.6$ are shifted by 200 and 400, respectively.
@@ III.1 Polar representation
We note that $\theta$ is simply used as the argument for the PDF of the angles $\alpha$, and that $\alpha$ is defined in the interval $(0,2\pi)$ because they refer to the polar angle of the position of the tracer (see Figure 4). Using the explicit form of matrix $\Sigma(t)$ given in eq. (7), we can write
@@ III.2 Turning angles
While the information given by the distribution of angles $\alpha_{t}$ is interesting, it is not necessarily useful in the case of real data, where the choice of any specific axis might be arbitrary. A more natural approach is to consider the so-called turning angles, generated by consecutive increments [59, 60, 61, 62, 63, 64, 65, 66]. In the correlated BM, the increments are still independent. This independence, as stated in the previous section, leads to the independence of the corresponding angles $\alpha_{t}^{\Delta}$. Specifically, let $\phi_{t}$ denote the angle between two consecutive increments at times $t$ and $t+1$. Figure 4 shows the angles of the increments $\alpha_{t}^{\Delta}$ and the turning angles $\phi_{t}$ in a sketched interval consisting of three consecutive observations within a trajectory. Angles $\alpha_{t}^{\Delta}$ are measured to the horizontal axis, preserving the counter-clockwise direction (cf. $\alpha_{2}^{\Delta}$ in the figure). The turning angles $\phi_{t}$ are defined by
@@
This definition of turning angles is directly connected to the velocity autocorrelation function, albeit without the ensemble average [60]. While in two dimensions it is possible to work with $\phi$ on the interval $[0,2\pi)$, the definition on the interval $[0,\pi)$, which arises from inverting eq. (13), can be directly extended to higher dimensions [65]. If a process involves torque or chirality, then the full circle should be considered but that is not the case here. For these reasons, most of the literature on turning angles currently adopts the interval $[0,\pi)$ [62, 63, 64, 65, 66] and we adhere to this convention. The connection to the angles $\alpha_{t}^{\Delta}$, described in the previous section, is as follows
$\phi_{t}\equiv\begin{cases}|\alpha_{t+1}^{\Delta}-\alpha_{t}^{\Delta}|,\quad&\mathrm{for}\ \alpha_{t+1}^{\Delta}-\alpha_{t}^{\Delta}\in(-\pi,\pi),\\ \\ 2\pi-(\alpha_{t+1}^{\Delta}-\alpha_{t}^{\Delta}),\quad\quad&\mathrm{for}\ \alpha_{t+1}^{\Delta}-\alpha_{t}^{\Delta}\in(\pi,2\pi),\\ \\ 2\pi+(\alpha_{t+1}^{\Delta}-\alpha_{t}^{\Delta}),\quad\quad&\mathrm{for}\ \alpha_{t+1}^{\Delta}-\alpha_{t}^{\Delta}\in(-2\pi,-\pi).\end{cases}$ (14)
From the properties of Brownian motion, the distribution of turning angles is symmetric about $\pi$, that is $g_{\phi}(\theta)=g_{\phi}(\pi-\theta)$ (this will also be verified later). Thus, we can employ a modified variable $\hat{\phi}$ that has the same distribution as $\phi$,
$\hat{\phi}_{t}\equiv\alpha_{t+1}^{\Delta}-\alpha_{t}^{\Delta}\ (\mathrm{mod}\ \pi).$ (15)
which makes it easier to analytically compute the distributions. In the rest of the article, we use $\hat{\phi}$ in all analytical derivations and $\phi$ in all numerical simulations. Both distributions are indeed found to be the same.
@@
For correlated BM, we can calculate the PDF of $\phi_{t}$ directly. Due to the increments of Brownian motion being independent and identically distributed (i.i.d.), we know that $\alpha_{t}^{\Delta}$’s are independent, and both $\alpha_{t}^{\Delta}$’s and $\phi_{t}$’s are identically distributed, and their distribution does not depend on time. This is not universally true and would not be the case for, for example, a two-dimensional fractional Brownian motion. Considering correlated BM, we can calculate the convolution of the PDF given in eq. (10). Then the density of $\alpha_{t+1}^{\Delta}-\alpha_{t}^{\Delta}$, prior to the modulo operation, is given by
$g_{{\hat{\phi}}}(\theta)$ $=(p_{\alpha}*\overline{p}_{-\alpha})(\theta)$
$=\frac{1-\rho^{2}}{4\pi^{2}}\int_{\max\{0,\theta\}}^{\min\{2\pi,2\pi+\theta\}}\frac{1}{\cos^{2}({x})\sqrt{\frac{D_{2}}{D_{1}}}+\sin^{2}({x})\sqrt{\frac{D_{1}}{D_{2}}}-2\rho\sin({x})\cos({x})}\times$
@@
$g_{{\hat{\phi}}}(\theta)$ $=(p_{\alpha}*\overline{p}_{-\alpha})(\theta)=\frac{1-\rho^{2}}{4\pi^{2}}\int_{\max\{0,\theta\}}^{\min\{2\pi,2\pi+\theta\}}\frac{1}{1-\rho\sin(2x)}\frac{1}{1+\rho\sin(2(\theta-x))}\mathrm{d}x,$ (17)
@@
If we again relax the assumption that $D_{1}=D_{2}$, for any $D_{1},D_{2}>0$, we take into consideration the modulo operation, we arrive at the final formula for the density of $\phi_{t}$
$p_{{\hat{\phi}}}(\theta)=g_{{\hat{\phi}}}(\theta)+g_{{\hat{\phi}}}(\theta+\pi)+g_{{\hat{\phi}}}(\theta-\pi)+g_{{\hat{\phi}}}(\theta-2\pi).$ (18)
Let us consider one more specific case, namely, uncorrelated coordinates with the same diffusion coefficient, i.e. the isotropic case. In such a case, we have
@@
$g_{{\hat{\phi}}}(\theta)=\begin{cases}\frac{1}{(2\pi)^{2}}\theta+\frac{1}{2\pi},\quad\theta\in(-2\pi,0),\\ \frac{-1}{(2\pi)^{2}}\theta+\frac{1}{2\pi},\quad\theta\in(0,2\pi).\end{cases}$ (19)
Having the formula for $g_{{\hat{\phi}}}$ in eq. (19), we can plug it into eq. (18) to obtain the PDF after modulo operation. Indeed, for $\theta\in(0,\pi)$ we have
$p_{{\hat{\phi}}}(\theta)$ $=g_{{\hat{\phi}}}(\theta)$ $+g_{{\hat{\phi}}}(\theta+\pi)$ $+g_{{\hat{\phi}}}(\theta-\pi)$ $+g_{{\hat{\phi}}}(\theta-2\pi)$
@@
For comparison, we consider $D_{1}=2,D_{2}=1$ and $\rho=0.7$. The PDF of the turning angles is presented in Figure 5. We can see a clear deviation from the uniform distribution. The solid blue histogram represents the turning angles derived from the simulated data (based on 1000 trajectories of length $2^{10}$ with $\delta=1$), while the dashed yellow line corresponds to eq. (18).
@@
Figure 5: Distribution of turning angles $\phi_{t}$ for the process with $D_{1}=2$, $D_{2}=1$, and $\rho=0.7$. The solid blue histogram corresponds to the angles calculated using eq. (13) from simulated data, and the dashed yellow line corresponds to $p_{\hat{\phi}}$ from eq. (18).
@@ IV Data
In this section, we present the results for two vastly different datasets. The first is a financial dataset that encompasses two of the main US market indices, the Dow Jones Industrial Average (DJIA) and the Standard and Poor’s 500 (S&P 500), for which we analyze their logarithmic returns. The second one relates to a soft-matter physical system, namely the motion of polystyrene beads in water. In Figure 6(a) we present the two-dimensional time series of the financial data considered, while in Figure 6(b) we demonstrate the example two-dimensional trajectory of polystyrene beads in water. Despite obtaining universal results for any diffusion coefficients $D_{1},D_{2}$ related to different directions, we first choose to normalize each of the coordinates considered by their corresponding standard deviation. While this normalization is not useful in physical systems such as the motion of a particle, it is natural in cases where a consistent unit system is not well defined, such as financial indices. In this case, normalization makes the system of units consistent. For a physical system, the normalization procedure is superfluous and is equivalent to non-dimensionalization.
[…]
Figure 6: Plots of representative sample trajectories of the analyzed datasets. (a) Financial data. (b) Beads in water.
@@ IV.1 Example: Financial data
In Figure 7 we present the analysis of the logarithmic returns of these two indices measured daily for a period of 2 years, from May 1, 2022, to May 1, 2024. Figure 7(a) shows the joint PDF, with the top and right histograms corresponding to DJIA and S&P500 logarithmic returns, respectively. In Figure 7(b), we present fitted Gaussian PDFs of the logarithmic returns of both indices over the selected period. The blue bars and dashed blue lines correspond to the DJIA index, while the yellow bars and dashed-dotted yellow lines correspond to S&P500. In Figure 7(c) we compare the empirical PDF of turning angles (blue bars) to the analytical density given by eq. (11). These results show a strong correlation between the two evaluated components, the DJIA and S&P500 logarithmic returns, as expected for two indices both focusing on U.S. stocks. These correlations serve to validate the presented approach where the turning angle distribution agrees well with our analytical predictions.
@@ IV.2 Example: Polystyrene beads in water
The data consists of 150 two-dimensional trajectories of polystyrene beads submersed in water imaged during $40.96$ s at a frame rate of $100$ frames/s [25, 67]. The nominal bead radius is $0.6\ \mu m$. The movement of the beads can be modeled using a classical Brownian motion without correlations between the coordinates ($\rho=0$). Due to the isotropic nature of the environment, we observe that $D_{1}$ and $D_{2}$ are indeed the same. The density of increments (presented in Figure 8(a)) has the circular contour lines of the characteristic of the joint PDF of two i.i.d. random variables. The one-dimensional histograms of the data are presented in Figure 8(b) and compared with the Gaussian PDFs with matched estimated parameters ($\langle D\rangle=0.375\mu\textrm{m}^{2}$/s). In this case, given $\rho=0$, the turning angles appear to have a uniform distribution (Figure 8(c)). In this case the characterization of the projections along the $x$ and $y$ directions, in terms of the turning angle distribution unequivocally shows the two components are not correlated, in agreement with the analytical predictions. The visual inspection of the 2D histogram also validates these conclusions.